OpenStudy (anonymous):

What are the x-intercept(s) of the graph of y − 8 = x2 − 6x?

OpenStudy (anonymous):

a. (−2, 0) and (4, 0) b. (2, 0) and (−4, 0) c. (2, 0) and (4, 0) d. (−2, 0) and (−4, 0)

OpenStudy (anonymous):

It's B right?

OpenStudy (ivettef365):

just substitute the y for 0, then you have when y=0 0 - 8 = x^2 -6x

OpenStudy (ivettef365):

then you have 0 = x^2 -6x + 8, do you know how to solve this

OpenStudy (ivettef365):

and b is not right

OpenStudy (anonymous):

I don't know how to solve it

OpenStudy (ivettef365):

0 = (x-2)(x-4) (when you multiply this you will get same equation x^2-6x+8) therefore you have x-2 = 0 and x - 4 = 0 x=2 and x=4, so the values would be (2,0) and (4,0)

OpenStudy (anonymous):

Thanks!